the scale on a map says that 2 inches equals 150 miles. If two cities are actually 325 miles apart, how many inches apart will they be on the map?
Answer by TimothyLamb(4379)

2/150 = i/325
i = (2/150)325 4.333 inches
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php